OpenStudy (anonymous):

A rectangle's length is 4 feet more than its width. If the area of the rectangle is 396 square feet, what is its width, in feet? ***please show step by step!! and explain!! thanks!! :D

OpenStudy (anonymous):

***No answer choices in this one :/

hartnn (hartnn):

isn't this same as last one ?

OpenStudy (anonymous):

its similar i think... but this time they give me the area.... so is the equation i use the one we found in the last one?

OpenStudy (unklerhaukus):

\[A_{\text{rectangle}}=\ell\times w\] \[\ell=(4+w)\]\[A_{\text{rectangle}}=396\]

hartnn (hartnn):

use that ^

OpenStudy (anonymous):

what @UnkleRhaukus wrote?

hartnn (hartnn):

Arectangle=ℓ×w ℓ=(4+w) Arectangle=396

OpenStudy (unklerhaukus):

\[396=(4+w)w\]

OpenStudy (anonymous):

so i get this? 396=4w+w^2

OpenStudy (unklerhaukus):

simplify the equation by bringing all the terms to one side, then solve, using the quadratic equation

OpenStudy (anonymous):